Cost of 4 dozen bananas is ₹ 180. How many bananas can be purchased for ₹ 90?

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Number of bananas bought in₹ 180= 4 dozens
= 4× 12
= 48 bananas
Number of bananas bought in ₹ 1 = 48 / 180
So, number of bananas bought in ₹90 = (48 / 180)×90
= 24 bananas
∴24 bananas can be purchased in₹ 90

Thanks